# Onsager reciprocal relations for charge and spin transport in periodically driven systems

### v1

@@ III.2 Bath Hamiltonian $H_{\textrm{b}}$
III.3 System-bath coupling Hamiltonian $H_{\textrm{sb}}$
$H_{\textrm{sb}}$ describes the coupling between the system and bath Tsuji ; Mikami ; NA-FloquetSHE :
@@ III.3 System-bath coupling Hamiltonian $H_{\textrm{sb}}$
where $V_{pa\sigma}$ is the coupling constant. In this study, we treat the effects of $H_{\textrm{sb}}$ as second-order perturbation. As a result, its main effect can be reduced to the damping Tsuji ; Mikami ; NA-FloquetSHE . Because of this damping, a nonequilibrium steady state could be achieved Tsuji ; Mikami ; NA-FloquetSHE .
IV Floquet linear-response theory
In this section, we formulate the time-averaged spin and charge off-diagonal dc conductivities in the nonequilibrium steady states of our periodically driven systems by using the Floquet linear-response theory Tsuji ; Mikami ; NA-FloquetSHE ; NA-Onsager ; NA-BCPL ; Eckstein . In this theory, we set $\mathbf{A}(t)=\mathbf{A}_{\textrm{pump}}(t)+\mathbf{A}_{\textrm{prob}}(t)$ and treat the effects of $\mathbf{A}_{\textrm{pump}}(t)$ in the Floquet theory Floquet1 ; Floquet2 and those of $\mathbf{A}_{\textrm{prob}}(t)$ in the linear-response theory Kubo . Therefore, this is a theory for the pump-probe measurements Opt-review in which the spin or charge current is generated perpendicular to the probe field with the pump field (Fig. 1). In addition, this theory is a natural extension of the Kubo formula Kubo for nondriven systems to the periodically driven systems.
IV.1 Time-averaged spin and charge off-diagonal dc conductivities $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$
Since the effects of $\mathbf{A}_{\textrm{prob}}(t)$ are treated in the linear-response theory Kubo , the spin and charge off-diagonal dc conductivities as functions of time, $\sigma_{yx}^{\textrm{S}}(t,t^{\prime})$ and $\sigma_{yx}^{\textrm{C}}(t,t^{\prime})$, are given by NA-FloquetSHE
@@ IV Floquet linear-response theory
where $\langle j_{\textrm{S}}^{y}(t)\rangle$ and $\langle j_{\textrm{C}}^{y}(t)\rangle$ are the expectation values of the operators of the spin current and charge current densities, respectively. These expectation values should be taken over the states with both the pump and probe fields. Then, the operators of the spin and charge currents are given by NA-FloquetSHE
@@
and $\textrm{sgn}(\sigma)=1$ or $-1$ for $\sigma=\uparrow$ or $\downarrow$, respectively. Note that Eq. (38) and (39) can be derived by using the continuity equations NA-FloquetSHE . We also note that $J_{\textrm{S}}^{y}(t)=Nj_{\textrm{S}}^{y}(t)$ and $J_{\textrm{C}}^{y}(t)=Nj_{\textrm{C}}^{y}(t)$ in our model due to $V=Na_{\textrm{NN}}^{2}=N$, where $V$ is the volume of the system.
$\sigma_{yx}^{\textrm{S}}(t,t^{\prime})$ and $\sigma_{yx}^{\textrm{C}}(t,t^{\prime})$ can be expressed in terms of nonequilibrium Green’s functions Eckstein ; NA-FloquetSHE . Substituting Eqs. (38) and (39) into Eqs. (36) and (37), respectively, we have
@@
where $\sigma_{yx}^{\textrm{Q}(1)}(t,t^{\prime})$ and $\sigma_{yx}^{\textrm{Q}(2)}(t,t^{\prime})$ (Q$=$S or C) are given by
@@
and $G_{b\sigma^{\prime}a\sigma}^{<}(\mathbf{k};t,t^{\prime})$ is the lesser Green’s function Mahan ; Keldysh ; Kadanoff-Baym ; NA-FloquetSHE ,
@@
We emphasize that the group velocities and Green’s function appearing in the right-hand sides of Eqs. (44) and (45) are those with only the pump field (i.e., no probe field). This is an advantage of the linear-response theory Kubo . In periodically driven cases, these quantities and the transport coefficients are of nonequilibrium. Meanwhile, in nondriven cases, these quantities are of equilibrium, although the transport coefficients are of nonequilibrium; this is sometimes misunderstood as if the transport coefficients were also of equilibrium. To rewrite Eq. (45), we use the equation,
$\frac{\delta G_{b\sigma a\sigma}^{<}(\mathbf{k};t,t)}{\delta A_{\textrm{prob}}^{x}(t^{\prime})}$
$=$ $-\sum_{c,d}\sum_{\sigma^{\prime}}v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k},t^{\prime})\Bigl{[}G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k};t,t^{\prime})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k};t^{\prime},t)$
@@
where $G_{a\sigma b\sigma^{\prime}}^{\textrm{R}}(\mathbf{k};t,t^{\prime})$ and $G_{a\sigma b\sigma^{\prime}}^{\textrm{A}}(\mathbf{k};t,t^{\prime})$ are the retarded and advanced Green’s functions Mahan ; Keldysh ; Kadanoff-Baym ; NA-FloquetSHE ,
@@
Note that Eq. (47) is obtained by using the Dyson equation of the Green’s functions and the Langreth rule Noneq-G-text ; Mikami ; Langreth . Substituting Eq. (47) into Eq. (45), we get
@@
×$\times\Bigl{[}G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k};t,t^{\prime})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k};t^{\prime},t)$
@@
The spin transport and charge transport in a nonequilibrium steady state of our periodically driven system can be described by the time-averaged spin and charge off-diagonal dc conductivities $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$, respectively [see Figs. 4(d) and 4(a)]. $\sigma_{yx}^{\textrm{Q}}$ (Q$=$S or C) is defined as
@@
[…]
$t_{\textrm{rel}}=t-t^{\prime}$, and $t_{\textrm{av}}=(t+t^{\prime})/2$. Note that the dc limit $\omega\rightarrow 0$ can be appropriately taken in the same way as for nondriven systems because the probe frequency $\omega$ is coupled to $t_{\textrm{rel}}$ in the Fourier transformation, whereas the time average is taken with respect to $t_{\textrm{av}}$. In Eq. (52), we have considered only the real part because we focus on the time-averaged dc conductivities in this paper. To calculate $\sigma_{yx}^{\textrm{Q}}(\omega)$, we use the Floquet representation of the Green’s functions Tsuji-PRB ; NA-FloquetSHE . In general, the Green’s functions in a periodically driven system depend not only on $t_{\textrm{rel}}$, but also on $t_{\textrm{av}}$. Because of this property, we should perform two transformations to convert the Green’s functions into the frequency functions:
[…]
×$\times G^{r}_{a\sigma b\sigma^{\prime}}(\mathbf{k};t_{\textrm{av}}+\frac{t_{\textrm{rel}}}{2},t_{\textrm{av}}-\frac{t_{\textrm{rel}}}{2}),$ (53)
where $G^{r}_{a\sigma b\sigma^{\prime};n}(\mathbf{k};\omega)$’s ($r=$ R, A, $<$) are defined in the range of $-\infty<\omega<\infty$. Since our Hamiltonian has the discrete time translation symmetry, we can restrict the range of frequency to $-\frac{\Omega}{2}\leq\omega<\frac{\Omega}{2}$. This property can be taken into account by performing another transformation:
@@
where $[G^{r}_{a\sigma b\sigma^{\prime}}(\mathbf{k},\omega)]_{mn}$’s ($-\frac{\Omega}{2}\leq\omega<\frac{\Omega}{2}$) are the Green’s functions in the Floquet representation Tsuji-PRB ; NA-FloquetSHE . By combining Eqs. (52), (42)–(44), and (50) and using the Floquet representation of the Green’s functions, we obtain
$\sigma_{yx}^{\textrm{Q}}(\omega)=\frac{1}{N}\sum_{\mathbf{k}}\sum_{a,b,c,d}\sum_{\sigma,\sigma^{\prime}}\int_{-\Omega/2}^{\Omega/2}$ $\frac{d\omega^{\prime}}{2\pi}\{\textrm{tr}[v_{ab\sigma}^{(\textrm{Q})y}(\mathbf{k})\frac{G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime}+\omega)-G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime}-\omega)}{2\omega}v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]$
$-\textrm{tr}[v_{ab\sigma}^{(\textrm{Q})y}(\mathbf{k})G_{b\sigma c\sigma^{\prime}}^{<}(\mathbf{k},\omega^{\prime})v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})\frac{G_{d\sigma^{\prime}a\sigma}^{\textrm{A}}(\mathbf{k},\omega^{\prime}+\omega)-G_{d\sigma^{\prime}a\sigma}^{\textrm{A}}(\mathbf{k},\omega^{\prime}-\omega)}{2\omega}]\},$ (55)
where $\textrm{tr}(ABCD)=\sum_{m,l,n,q=-\infty}^{\infty}[A]_{ml}[B]_{ln}[C]_{nq}[D]_{qm}$, and $m$, $l$, $n$, and $q$ are indices in the Floquet representation. For the derivation of Eq. (55), see Appendix A. Note that $[v_{ab\sigma}^{(\textrm{Q})\nu}(\mathbf{k})]_{mn}$ ($\nu=x,y$) is given by
@@
×$\times\{\textrm{tr}[v_{ab\sigma}^{(\textrm{S})y}(\mathbf{k})\frac{\partial G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime})}{\partial\omega^{\prime}}v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]$
@@
×$\times\{\textrm{tr}[v_{ab\sigma}^{(\textrm{C})y}(\mathbf{k})\frac{\partial G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime})}{\partial\omega^{\prime}}v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]$
@@
The Green’s functions appearing in Eqs. (57) and (58) can be determined in the following way. For our periodically driven system, the Green’s functions are determined from the Dyson equation in the matrix form Keldysh ; Keldysh2 ; NA-FloquetSHE ,
@@
where $\Gamma$ is the damping. In deriving these equations, we have omitted the real parts and replaced $\pi\sum_{p}V_{pa\sigma}V_{pb\sigma^{\prime}}\delta(\omega+m\Omega-\epsilon_{p}+\mu_{\textrm{b}})$ by $\Gamma\delta_{a,b}\delta_{\sigma,\sigma^{\prime}}$ for simplicity NA-FloquetSHE . Such simplification may be sufficient because the main effect of $H_{\textrm{sb}}$ is to induce the damping, which makes the system the nonequilibrium steady state Tsuji ; Mikami ; NA-FloquetSHE . Then, Eq. (59) can be rewritten as
@@
For $[\epsilon_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL, see Appendix B; for that of Sr2RuO4 driven by CPL or LPL, see Ref. [NA-FloquetSHE, ]. In deriving Eq. (82), we have chosen the Keldysh component of the matrix $G_{0}^{-1}$ to be zero because it contains the information about the initial condition. This treatment may be valid to describe a nonequilibrium steady state with finite damping Oka-PRB ; Mikami ; Tsuji ; NA-FloquetSHE because a nonequilibrium steady state should be independent of a choice of the initial condition. Note that the retarded and advanced components of the matrix $G_{0}^{-1}$ are obtained by replacing $\Gamma$ in Eqs. (80) and (81), respectively, by a positive infinitesimal. Because of the matrix relation $G^{-1}G=1$, the components of the inverse matrices satisfy
@@
IV.2 The other off-diagonal dc conductivities $\tilde{\sigma}_{xy}^{\textrm{S}}$, $\sigma_{xy}^{\textrm{C}}$, $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$, and $\bar{\sigma}_{xy}^{\textrm{C}}$
To discuss the Onsager reciprocal relations, we need to consider $\tilde{\sigma}_{xy}^{\textrm{S}}$ and $\sigma_{xy}^{\textrm{C}}$. Here $\tilde{\sigma}_{xy}^{\textrm{S}}$ or $\sigma_{xy}^{\textrm{C}}$ is the time-averaged off-diagonal dc conductivity for the charge current along the $x$ axis perpendicular to the probe spin or electric field, respectively, applied along the $y$ axis with the pump field $\mathbf{A}_{\textrm{pump}}(t)$ [see Figs. 4(f) and 4(c)]. These conductivities are given by
@@
×$\times\{\textrm{tr}[v_{ab\sigma}^{(\textrm{C})x}(\mathbf{k})\frac{\partial G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime})}{\partial\omega^{\prime}}v_{cd\sigma^{\prime}}^{(\textrm{S})y}(\mathbf{k})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]$
@@
×$\times\{\textrm{tr}[v_{ab\sigma}^{(\textrm{C})x}(\mathbf{k})\frac{\partial G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime})}{\partial\omega^{\prime}}v_{cd\sigma^{\prime}}^{(\textrm{C})y}(\mathbf{k})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]$
@@
In addition, we need to consider $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$ and $\bar{\sigma}_{xy}^{\textrm{C}}$. Here $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$ or $\bar{\sigma}_{xy}^{\textrm{C}}$ is the time-averaged off-diagonal dc conductivity for the charge current along the $x$ axis perpendicular to the probe spin or electric field, respectively, applied along the $y$ axis with the pump field $\mathbf{A}_{\textrm{pump}}(-t)$ [see Figs. 4(e) and 4(b)]. These conductivities are given by
@@
×$\times\{\textrm{tr}[\bar{v}_{ab\sigma}^{(\textrm{C})x}(\mathbf{k})\frac{\partial\bar{G}_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime})}{\partial\omega^{\prime}}\bar{v}_{cd\sigma^{\prime}}^{(\textrm{S})y}(\mathbf{k})\bar{G}_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]$
@@
×$\times\{\textrm{tr}[\bar{v}_{ab\sigma}^{(\textrm{C})x}(\mathbf{k})\frac{\partial\bar{G}_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime})}{\partial\omega^{\prime}}\bar{v}_{cd\sigma^{\prime}}^{(\textrm{C})y}(\mathbf{k})\bar{G}_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]$
@@
Here $[\bar{v}_{ab\sigma}^{(\textrm{Q})\nu}(\mathbf{k})]_{mn}$ ($\nu=x,y$) is given by
@@
and $[\bar{G}^{r}_{a\sigma b\sigma^{\prime}}(\mathbf{k},\omega)]_{mn}$’s are determined by replacing Eqs. (80)–(86) by

### v2

@@ 3.2 Bath Hamiltonian $H_{\textrm{b}}$
3.3 System-bath coupling Hamiltonian $H_{\textrm{sb}}$
$H_{\textrm{sb}}$ describes the coupling between the system and bath [38, 8, 11]:
@@ 3.3 System-bath coupling Hamiltonian $H_{\textrm{sb}}$
where $V_{pa\sigma}$ is the coupling constant. In this study, we treat the effects of $H_{\textrm{sb}}$ as second-order perturbation. As a result, its main effect can be reduced to the damping [38, 8, 11]. Because of this damping, a nonequilibrium steady state could be achieved [38, 8, 11].
4 Floquet linear-response theory
In this section, we formulate the time-averaged spin and charge off-diagonal dc conductivities in the nonequilibrium steady states of our periodically driven systems by using the Floquet linear-response theory [38, 8, 11, 27, 30, 49]. In this theory, we set $\mathbf{A}(t)=\mathbf{A}_{\textrm{pump}}(t)+\mathbf{A}_{\textrm{prob}}(t)$ and treat the effects of $\mathbf{A}_{\textrm{pump}}(t)$ in the Floquet theory [1, 2] and those of $\mathbf{A}_{\textrm{prob}}(t)$ in the linear-response theory [21]. Therefore, this is a theory for the pump-probe measurements [50] in which the spin or charge current is generated perpendicular to the probe field with the pump field (Fig. 1). In addition, this theory is a natural extension of the Kubo formula [21] for nondriven systems to the periodically driven systems.
4.1 Time-averaged spin and charge off-diagonal dc conductivities $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$
Since the effects of $\mathbf{A}_{\textrm{prob}}(t)$ are treated in the linear-response theory [21], the spin and charge off-diagonal dc conductivities as functions of time, $\sigma_{yx}^{\textrm{S}}(t,t^{\prime})$ and $\sigma_{yx}^{\textrm{C}}(t,t^{\prime})$, are given by [11]
@@ 4 Floquet linear-response theory
where $\langle j_{\textrm{S}}^{y}(t)\rangle$ and $\langle j_{\textrm{C}}^{y}(t)\rangle$ are the expectation values of the operators of the spin current and charge current densities, respectively. These expectation values should be taken over the states with both the pump and probe fields. Then, the operators of the spin and charge currents are given by [11]
@@
and $\textrm{sgn}(\sigma)=1$ or 1-1 for $\sigma=\uparrow$ or $\downarrow$, respectively. Note that Eq. (38) and (39) can be derived by using the continuity equations [11]. We also note that $J_{\textrm{S}}^{y}(t)=Nj_{\textrm{S}}^{y}(t)$ and $J_{\textrm{C}}^{y}(t)=Nj_{\textrm{C}}^{y}(t)$ in our model due to $V=Na_{\textrm{NN}}^{2}=N$, where $V$ is the volume of the system.
$\sigma_{yx}^{\textrm{S}}(t,t^{\prime})$ and $\sigma_{yx}^{\textrm{C}}(t,t^{\prime})$ can be expressed in terms of nonequilibrium Green’s functions [49, 11]. Substituting Eqs. (38) and (39) into Eqs. (36) and (37), respectively, we have
@@
where $\sigma_{yx}^{\textrm{Q}(1)}(t,t^{\prime})$ and $\sigma_{yx}^{\textrm{Q}(2)}(t,t^{\prime})$ (Q==S or C) are given by
@@
and $G_{b\sigma^{\prime}a\sigma}^{<}(\mathbf{k};t,t^{\prime})$ is the lesser Green’s function [51, 52, 53, 11],
@@
We emphasize that the group velocities and Green’s function appearing in the right-hand sides of Eqs. (44) and (45) are those with only the pump field (i.e., no probe field). This is an advantage of the linear-response theory [21]. In periodically driven cases, these quantities and the transport coefficients are of nonequilibrium. Meanwhile, in nondriven cases, these quantities are of equilibrium, although the transport coefficients are of nonequilibrium; this is sometimes misunderstood as if the transport coefficients were also of equilibrium. To rewrite Eq. (45), we use the equation,
$\frac{\delta G_{b\sigma a\sigma}^{<}(\mathbf{k};t,t)}{\delta A_{\textrm{prob}}^{x}(t^{\prime})}$ $=-\sum_{c,d}\sum_{\sigma^{\prime}}v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k},t^{\prime})$
$\times\Bigl{[}G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k};t,t^{\prime})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k};t^{\prime},t)$
@@
where $G_{a\sigma b\sigma^{\prime}}^{\textrm{R}}(\mathbf{k};t,t^{\prime})$ and $G_{a\sigma b\sigma^{\prime}}^{\textrm{A}}(\mathbf{k};t,t^{\prime})$ are the retarded and advanced Green’s functions [51, 52, 53, 11],
@@
Note that Eq. (47) is obtained by using the Dyson equation of the Green’s functions and the Langreth rule [54, 8, 55]. Substituting Eq. (47) into Eq. (45), we get
@@
$\times\Bigl{[}G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k};t,t^{\prime})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k};t^{\prime},t)$
@@
The spin transport and charge transport in a nonequilibrium steady state of our periodically driven system can be described by the time-averaged spin and charge off-diagonal dc conductivities $\sigma_{yx}^{\textrm{S}}$ and $\sigma_{yx}^{\textrm{C}}$, respectively [see Figs. 4(d) and 4(a)]. $\sigma_{yx}^{\textrm{Q}}$ (Q==S or C) is defined as
@@
$\sigma_{yx}^{\textrm{Q}}(\omega)=\textrm{Re}\int_{0}^{T_{\textrm{p}}}\frac{dt_{\textrm{av}}}{T_{\textrm{p}}}\int_{-\infty}^{\infty}dt_{\textrm{rel}}e^{i\omega t_{\textrm{rel}}}\sigma_{yx}^{\textrm{Q}}(t,t^{\prime}),$ (52)
$t_{\textrm{rel}}=t-t^{\prime}$, and $t_{\textrm{av}}=(t+t^{\prime})/2$. Note that the dc limit $\omega\rightarrow 0$ can be appropriately taken in the same way as for nondriven systems because the probe frequency $\omega$ is coupled to $t_{\textrm{rel}}$ in the Fourier transformation, whereas the time average is taken with respect to $t_{\textrm{av}}$. In Eq. (52), we have considered only the real part because we focus on the time-averaged dc conductivities in this paper. To calculate $\sigma_{yx}^{\textrm{Q}}(\omega)$, we use the Floquet representation of the Green’s functions [56, 11]. In general, the Green’s functions in a periodically driven system depend not only on $t_{\textrm{rel}}$, but also on $t_{\textrm{av}}$. Because of this property, we should perform two transformations to convert the Green’s functions into the frequency functions:
$G^{r}_{a\sigma b\sigma^{\prime};n}(\mathbf{k};\omega)$ $=\int_{-\infty}^{\infty}dt_{\textrm{rel}}e^{i\omega t_{\textrm{rel}}}\int_{0}^{T_{\textrm{p}}}\frac{dt_{\textrm{av}}}{T_{\textrm{p}}}e^{in\Omega t_{\textrm{av}}}$
$\times G^{r}_{a\sigma b\sigma^{\prime}}(\mathbf{k};t_{\textrm{av}}+\frac{t_{\textrm{rel}}}{2},t_{\textrm{av}}-\frac{t_{\textrm{rel}}}{2}),$ (53)
where $G^{r}_{a\sigma b\sigma^{\prime};n}(\mathbf{k};\omega)$’s ($r=$ R, A, <<) are defined in the range of $\infty<\omega<\infty$. Since our Hamiltonian has the discrete time translation symmetry, we can restrict the range of frequency to $\frac{\Omega}{2}\leq\omega<\frac{\Omega}{2}$. This property can be taken into account by performing another transformation:
@@
where $G^{r}_{a\sigma b\sigma^{\prime}}(\mathbf{k},\omega)]_{mn}$’s ($\frac{\Omega}{2}\leq\omega<\frac{\Omega}{2}$) are the Green’s functions in the Floquet representation [56, 11]. By combining Eqs. (52), (42)–(44), and (50) and using the Floquet representation of the Green’s functions, we obtain
$\sigma_{yx}^{\textrm{Q}}(\omega)=\frac{1}{N}\sum_{\mathbf{k}}\sum_{a,b,c,d}\sum_{\sigma,\sigma^{\prime}}\int_{-\Omega/2}^{\Omega/2}\frac{d\omega^{\prime}}{2\pi}$
$\times\{\textrm{tr}[v_{ab\sigma}^{(\textrm{Q})y}(\mathbf{k})\frac{G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime}+\omega)-G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime}-\omega)}{2\omega}$
$\ \ \ \ \ \ \times v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]$
$-\textrm{tr}[v_{ab\sigma}^{(\textrm{Q})y}(\mathbf{k})G_{b\sigma c\sigma^{\prime}}^{<}(\mathbf{k},\omega^{\prime})v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})$
$\ \ \ \ \ \times\frac{G_{d\sigma^{\prime}a\sigma}^{\textrm{A}}(\mathbf{k},\omega^{\prime}+\omega)-G_{d\sigma^{\prime}a\sigma}^{\textrm{A}}(\mathbf{k},\omega^{\prime}-\omega)}{2\omega}]\},$ (55)
where $\textrm{tr}(ABCD)=\sum_{m,l,n,q=-\infty}^{\infty}[A]_{ml}[B]_{ln}[C]_{nq}[D]_{qm}$, and $m$, $l$, $n$, and $q$ are indices in the Floquet representation. For the derivation of Eq. (55), see Appendix A. Note that $v_{ab\sigma}^{(\textrm{Q})\nu}(\mathbf{k})]_{mn}$ ($\nu=x,y$) is given by
@@
$\times\{\textrm{tr}[v_{ab\sigma}^{(\textrm{S})y}(\mathbf{k})\frac{\partial G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime})}{\partial\omega^{\prime}}v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]$
@@
$\times\{\textrm{tr}[v_{ab\sigma}^{(\textrm{C})y}(\mathbf{k})\frac{\partial G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime})}{\partial\omega^{\prime}}v_{cd\sigma^{\prime}}^{(\textrm{C})x}(\mathbf{k})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]$
@@
The Green’s functions appearing in Eqs. (57) and (58) can be determined in the following way. For our periodically driven system, the Green’s functions are determined from the Dyson equation in the matrix form [52, 57, 11],
@@
where $\Gamma$ is the damping. In deriving these equations, we have omitted the real parts and replaced $\pi\sum_{p}V_{pa\sigma}V_{pb\sigma^{\prime}}\delta(\omega+m\Omega-\epsilon_{p}+\mu_{\textrm{b}})$ by $\Gamma\delta_{a,b}\delta_{\sigma,\sigma^{\prime}}$ for simplicity [11]. Such simplification may be sufficient because the main effect of $H_{\textrm{sb}}$ is to induce the damping, which makes the system the nonequilibrium steady state [38, 8, 11]. Then, Eq. (59) can be rewritten as
@@
For $\epsilon_{ab}(\mathbf{k})]_{mn}$ of Sr2RuO4 driven by BCPL, see Appendix B; for that of Sr2RuO4 driven by CPL or LPL, see Ref. [[11]]. In deriving Eq. (82), we have chosen the Keldysh component of the matrix $G_{0}^{-1}$ to be zero because it contains the information about the initial condition. This treatment may be valid to describe a nonequilibrium steady state with finite damping [7, 8, 38, 11] because a nonequilibrium steady state should be independent of a choice of the initial condition. Note that the retarded and advanced components of the matrix $G_{0}^{-1}$ are obtained by replacing $\Gamma$ in Eqs. (80) and (81), respectively, by a positive infinitesimal. Because of the matrix relation $G^{-1}G=1$, the components of the inverse matrices satisfy
@@
4.2 The other off-diagonal dc conductivities $\tilde{\sigma}_{xy}^{\textrm{S}}$, $\sigma_{xy}^{\textrm{C}}$, $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$, and $\bar{\sigma}_{xy}^{\textrm{C}}$
To discuss the Onsager reciprocal relations, we need to consider $\tilde{\sigma}_{xy}^{\textrm{S}}$ and $\sigma_{xy}^{\textrm{C}}$. Here $\tilde{\sigma}_{xy}^{\textrm{S}}$ or $\sigma_{xy}^{\textrm{C}}$ is the time-averaged off-diagonal dc conductivity for the charge current along the $x$ axis perpendicular to the probe spin field or electric field, respectively, applied along the $y$ axis with the pump field $\mathbf{A}_{\textrm{pump}}(t)$ [see Figs. 4(f) and 4(c)]. These conductivities are given by
@@
$\times\{\textrm{tr}[v_{ab\sigma}^{(\textrm{C})x}(\mathbf{k})\frac{\partial G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime})}{\partial\omega^{\prime}}v_{cd\sigma^{\prime}}^{(\textrm{S})y}(\mathbf{k})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]$
@@
$\times\{\textrm{tr}[v_{ab\sigma}^{(\textrm{C})x}(\mathbf{k})\frac{\partial G_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime})}{\partial\omega^{\prime}}v_{cd\sigma^{\prime}}^{(\textrm{C})y}(\mathbf{k})G_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]$
@@
In addition, we need to consider $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$ and $\bar{\sigma}_{xy}^{\textrm{C}}$. Here $\bar{\tilde{\sigma}}_{xy}^{\textrm{S}}$ or $\bar{\sigma}_{xy}^{\textrm{C}}$ is the time-averaged off-diagonal dc conductivity for the charge current along the $x$ axis perpendicular to the probe spin field or electric field, respectively, applied along the $y$ axis with the pump field $\mathbf{A}_{\textrm{pump}}(-t)$ [see Figs. 4(e) and 4(b)]. These conductivities are given by
@@
$\times\{\textrm{tr}[\bar{v}_{ab\sigma}^{(\textrm{C})x}(\mathbf{k})\frac{\partial\bar{G}_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime})}{\partial\omega^{\prime}}\bar{v}_{cd\sigma^{\prime}}^{(\textrm{S})y}(\mathbf{k})\bar{G}_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]$
@@
$\times\{\textrm{tr}[\bar{v}_{ab\sigma}^{(\textrm{C})x}(\mathbf{k})\frac{\partial\bar{G}_{b\sigma c\sigma^{\prime}}^{\textrm{R}}(\mathbf{k},\omega^{\prime})}{\partial\omega^{\prime}}\bar{v}_{cd\sigma^{\prime}}^{(\textrm{C})y}(\mathbf{k})\bar{G}_{d\sigma^{\prime}a\sigma}^{<}(\mathbf{k},\omega^{\prime})]$
@@
Here $\bar{v}_{ab\sigma}^{(\textrm{Q})\nu}(\mathbf{k})]_{mn}$ ($\nu=x,y$) is given by
@@
and $\bar{G}^{r}_{a\sigma b\sigma^{\prime}}(\mathbf{k},\omega)]_{mn}$’s are determined by replacing Eqs. (80)–(86) by